### v1

# Single-Source Regular Path Querying in Terms of Linear Algebra

Abstract.
A given edge-labelled graph two-way regular path queries (2-RPQs) allow one to use regular languages over labelled edges and inverted edges to constraint paths of interest. 2-RPQs are (partially) adopted in different real-world graph analysis systems and are a part of the GQL ISO standard. But the performance of 2-RPQs on real-world graphs is still a bottleneck for wider adoption. A new single-source 2-RPQ algorithm based on linear algebra is proposed. Utilization of high-performance sparse linear algebra libraries for the algorithm implementation allows one to achieve significant speedup over competitors on real-world data and queries. Our implementation demonstrates better performance on average on Wikidata and the respective query log in comparison with MillenniumDB, FalkorDB, and the algorithm of Diego Arroyuelo et al.
Graph Querying, Regular Path Querying, RPQ, Graph Analysis Algorithms, Sparse Linear Algebra, GraphBLAS
copyright: acmlicensedjournalyear: 2024conference: ; ; booktitle: ccs: Information systems Graph-based database modelsccs: Information systems Query languages for non-relational enginesccs: Theory of computation Regular languagesccs: Mathematics of computing Paths and connectivity problems
1. Introduction
Language-constrained path querying (Barrett et al., 2000) is a way to search for paths in edge-labeled graphs where constraints are formulated in terms of a formal language. The language restricts the set of accepted paths: the sentence formed by the labels of a path should be in the language. Regular languages are the most popular class of constraints used as navigational queries in graph databases.
Queries that use regular languages to specify constraints are called regular path queries or RPQs. Being proposed in 1989 by Alberto O. Mendelzon and Peter T. Wood (Mendelzon and Wood, 1989), RPQs studying intensively for now. Different extensions of RPQs are studying too: two-way RPQ or 2-RPQ that extend the alphabet of RPQs by the inverse of relationship symbols (Calvanese et al., 2000a), conjunctive RPQ or CRPQ that allows one to check several (parallel) paths (Calvanese et al., 2000b). Regular path constraints and related extensions adopted (partially, in some cases) in a wide range of graph analysis systems and respective query languages, such as Cypher (Francis et al., 2018) and PGQL (van Rest et al., 2016). Moreover, RPQ is a part of ISO standard for a graph query language GQL [^1] ISO/IEC 39075:2024, GQL, https://www.iso.org/standard/76120.html, and core of SPARQL 1.1 RDF query language (201, 2013).
Nevertheless, having a long history of theoretical and applied research, adoption in real-world solutions, RPQs (and related extensions) are still in the focus of research. One of the important direction is an implementation and optimization of RPQ evaluation algorithms (Bonifati et al., 2018) to achieve better performance in real-world cases, and performance-targeted solutions still actively developed (Arroyuelo et al., 2022, 2023b; Ma et al., 2024). Thus, the development of new performant algorithms for RPQs evaluation is an actual challenge that also pointed out by Angela Bonifati et al in (Bonifati et al., 2018). Set of different solutions was proposed to improve RPQ evaluation performance: from special indexes creation (Liu et al., 2021; Arroyuelo et al., 2023b) to utilization of different parallel and distributed models (Nolé and Sartiani, 2016; Faltín et al., 2023; Guo et al., 2021a; Wang et al., 2019).
Sparse linear algebra is a promising way to high-performance graph analysis. This idea is actively developed by GraphBLAS (Kepner et al., 2016) community [^2] GraphBLAS community web page: https://graphblas.org/. Huge amount of graph analysis algorithms, such as PageRank or triangle centralities, can be expressed in terms of linear algebra [^3] Almost full list of graph analysis algorithms in terms of linear algebra: https://github.com/GraphBLAS/GraphBLAS-Pointers. and respective implementations demonstrate promising performance in real-world cases (Pelletier et al., 2021). Even more recently, it was shown that utilization of sparse linear algebra is a way to a high-performance algorithm for more expressive class of queries, Context-Free Path Queries, or CFPQ (Terekhov et al., 2021). But, to our knowledge, there are only a few results on linear algebra based algorithms for RPQ. On the one hand, modern graph database FalkorBD444 Sources of FalkorBD on GitHub: https://github.com/FalkorDB/falkordb (formerly RedisGraph) (Cailliau et al., 2019) is based on SuiteSparse:GraphBLAS (Davis, 2023), reference implementation of the GraphBLAS API, and utilises linear algebra to analyse graphs. This database supports a subset of the Cypher graph query language, including parts of regular constraints. But there is no detailed analysis of respective algorithm. At the same time, linear algebra based RPQ evaluation algorithm was recently proposed by Diego Arroyuelo et al (Arroyuelo et al., 2023a). The proposed solution directly translates regular expressions to expressions over matrices and vectors. Being implemented using sparse matrices and parallel computations, this solution still has performance problems that are evident from the evaluation (Arroyuelo et al., 2023a).
There are BFS-based strategies for RPQ evaluation (e.g. (Koschmieder and Leser, 2012)) [^5] For overview of the primary techniques to evaluate regular path queries, such as relational algebra based and automata based approaches, and ways to optimise it, we refer to the book “Querying Graphs” by Angela Bonifati et al (Bonifati et al., 2018).. At the same time, BFS and its variations can be expressed in terms of linear algebra (Davis, 2019a). It opens a way to develop a linear algebra based algorithm for RPQ evaluation having BFS-like traversal as the core, and we investigate this way.
To summarise, in this work we make the following contributions.
  1. (1)
    A new BFS-inspired algorithm (LA 2-RPQ) for single-source (or single-destination as a symmetric case) 2-RPQ is proposed. The algorithm is based on linear algebra: it is expressed in terms of operations over sparse boolean matrices. This fact allows one to utilise high-performance parallel libraries, such as SuiteSparse:GraphBLAS (Davis, 2023), for 2-RPQ evaluation. The correctness of the proposed algorithm is proven.
  2. (2)
    Our implementation of the proposed algorithm, based on SuiteSparse:GraphBLAS, is evaluated and compared with other linear algebra based solutions such as FalkorDB and the algorithm proposed by Diego Arroyuelo et al (Arroyuelo et al., 2023a), and with the state-of-the-art graph database MillenniumDB. As a result, we show that our solution is comparable with competitors on real-world dataset Wikidata and the respective query log proposed in MillenniumDB path query challenge (Farias et al., 2023). While for some queries the proposed algorithm is slower than the competitors, it is faster on average: ×$\times 3.9$ for MillenniumDB, ×$\times 19.6$ for FalkorDB, ×$\times 4.2$ for the algorithm of Diego Arroyuelo et al. Moreover, it fits within the 1 minute time limit for all queries, while the other competitors require more time for some queries.
2. Preliminaries
In this section, we provide the theoretical basics of graph theory and formal language theory required to define the RPQ problem and to describe our solution.
2.1. Basics of graph theory
First we define the graph that we use as a data model.
Definition 2.0 (Edge-labelled graph).
A quadruple
$\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ is called an edge-labelled graph (or a graph) if:
  • $V$ is a finite set of vertices;
  • $E\subseteq V\times V$ is a finite set of edges;
  • $L$ is a finite set of labels;
  • $\lambda_{\mathcal{G}}:E\mapsto 2^{L}$ represents edge labels.
Any finite set can be enumerated by natural numbers from $1$ to $n$. For the rest of the paper, we will assume that the vertices of the graph $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ are enumerated, and without loss of generality $V=\{1,2,...,|V|\}$.
Let $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ be an edge-labelled graph. Introduce symbols and sets:
  • $a^{-}\notin L$ for $a\in L$, $(a^{-})^{-}=a$ and $a=b\Leftrightarrow a^{-}=b^{-}$
  • $e^{-}=(v,u)$ where $e=(u,v)$.
  • $E^{-}=\{e^{-}\mid e\in E\}$.
  • $L^{-}=\{a^{-}\mid a\in L\}$.
  • $\lambda_{\mathcal{G}}^{-}:E^{-}\mapsto 2^{L^{-}},\;\lambda_{\mathcal{G}}^{-}(e^{-})=\{a^{-}\mid a\in\lambda_{\mathcal{G}}(e)\}$.
We also need these sets to generalize the definition of the directed graph to be able to traverse it in both directions:
  • $E^{\leftrightarrow}=E\cup E^{-}$.
  • $L^{\leftrightarrow}=L\cup L^{-}$.
  • $\lambda_{\mathcal{G}}^{\leftrightarrow}:E^{\leftrightarrow}\mapsto 2^{L^{\leftrightarrow}},\;\;\lambda_{\mathcal{G}}^{\leftrightarrow}(e)=\lambda_{\mathcal{G}}(e)\cup\lambda^{-}_{\mathcal{G}}(e)\;$.
  • $\mathcal{G}^{\leftrightarrow}=\langle V,E^{\leftrightarrow},L^{\leftrightarrow},\lambda^{\leftrightarrow}\rangle$.
Definition 2.0 (path).
A path $\pi=(e_{1},e_{2},...,e_{n})$ of length $n$ in the graph $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ from $u_{1}$ to $v_{n}$ is a finite sequence of edges $e_{i}=(u_{i},v_{i})\in E$ s.t. $\forall 1\leq i\leq n-1\;v_{i}=u_{i+1}$.
We say there are zero-length paths represented by an empty sequence from $v$ to $v$ for all $v\in V$.
Definition 2.0 (2-way path).
A 2-way path (2-path) $\pi=(e_{1},e_{2},...,e_{n})$ in the graph $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ is a path in $\mathcal{G}^{\leftrightarrow}=\langle V,E^{\leftrightarrow},L^{\leftrightarrow},\lambda_{\mathcal{G}}^{\leftrightarrow}\rangle$.
The $\omega$ maps paths to words as defined below.
  • $\omega_{\mathcal{G}}(\pi)=\{a_{1}\cdot a_{2}\cdot...\cdot a_{n}\mid a_{i}\in\lambda_{\mathcal{G}}(e_{i})\}$ for path $\pi=(e_{1},e_{2},...,e_{n})$ in $\mathcal{G}$.
  • $\omega^{\leftrightarrow}_{\mathcal{G}}(\pi)=\{a_{1}\cdot a_{2}\cdot...\cdot a_{n}\mid a_{i}\in\lambda^{\leftrightarrow}_{\mathcal{G}}(e_{i})\}$ for 2-path $\pi=(e_{1},e_{2},...,e_{n})$ in $\mathcal{G}$.
where $\cdot$ denotes concatenation.
Regular languages (RLs) represent the set of all languages that are accepted by finite automata. However, we are going to look for 2-way paths for which we need an NFA modification to be introduced.
Definition 2.0 (2-way non-deterministic finite automaton).
A 2-way non-deterministic finite automaton (2-NFA) is a tuple $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$, where:
  • $Q$ is a finite set of states;
  • $\Sigma$ is a finite alphabet;
  • $\Delta\subseteq Q\times Q$ is transition relation;
  • $\lambda_{\mathcal{N}}:\Delta\mapsto 2^{\Sigma^{\leftrightarrow}}$ maps transitions to labels;
  • $Q_{S}\subseteq Q$ is a set of starting states;
  • $Q_{F}\subseteq Q$ is a set of final states.
The set of languages accepted by 2-NFAs coincides with the set of all RLs over $\Sigma^{\leftrightarrow}$. Let $[\![\mathcal{N}]\!]=\mathcal{R}$ where $\mathcal{R}$ is the RL accepted by the 2-NFA $\mathcal{N}$.
In practice special kinds of strings called regular expressions provide a useful facility to define regular languages. The algorithms supplying the transformations between regular expressions, RLs and 2-NFAs are out of the scope of this paper.
Notice that 2-NFA can be seen as a graph $\mathcal{G}_{\mathcal{N}}=\langle Q,\Delta,\Sigma^{\leftrightarrow},\lambda_{\mathcal{N}}\rangle$ equipped with the set of starting states $Q_{S}\subseteq Q$ and the set of final states $Q_{F}\subseteq Q$. Analogously introduce sets $\Delta^{-}$, $\lambda_{\mathcal{N}}^{-}$ and the map $\omega_{\mathcal{N}}=\{a_{1}\cdot a_{2}\cdot...\cdot a_{n}\mid a_{i}\in\lambda_{\mathcal{N}}(\delta_{i})\}$ for the paths $\pi=(\delta_{1},\delta_{2},...,\delta_{n})$ in the graph $\mathcal{G}_{\mathcal{N}}$.
Vice versa a graph $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ equipped with some $V_{S}\subseteq V$ can bee seen as 2-NFA $\mathcal{N}_{\mathcal{G}}=\langle V,L,E,\lambda_{\mathcal{G}}^{\leftrightarrow},V_{S},V\rangle$. Such 2-NFA accepts the language:
$[\![\mathcal{N}_{G}]\!]=\bigcup_{\begin{subarray}{c}\text{2-path }\pi_{\mathcal{G}}\text{ in }\mathcal{G}\\ \text{ from }v_{S}\in V_{S}\end{subarray}}\omega_{\mathcal{G}}^{\leftrightarrow}(\pi_{\mathcal{G}})$
Thus, we can treat evaluation of 2-RPQs with fixed set of starting vertices as an intersection of 2-NFAs.
Definition 2.0 (2-NFA intersection).
For two arbitrary 2-NFAs $\mathcal{N}_{1}=\langle Q_{1},\Sigma_{1},\Delta_{1},\lambda_{\mathcal{N}_{1}},Q_{S1},Q_{F1}\rangle$ and $\mathcal{N}_{2}=\langle Q_{2},\Sigma_{2},\Delta_{2},\lambda_{\mathcal{N}_{2}},Q_{S2},Q_{F2}\rangle$ introduce a new automaton $\mathcal{N}=\mathcal{N}_{1}\times\mathcal{N}_{2}$ called the intersection of 2-NFAs $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ where $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$ s.t.:
  • $Q=Q_{1}\times Q_{2}$;
  • $\Sigma=\Sigma_{1}\cap\Sigma_{2}$;
  • $\Delta=\{((q_{1},q_{2}),(q_{1}^{\prime},q_{2}^{\prime}))\mid(q_{1},q_{1}^{\prime})\in\Delta_{1}$, $(q_{2},q_{2}^{\prime})\in\Delta_{2}\}$;
  • $\lambda_{\mathcal{N}}(((q_{1},q_{2}),(q_{1}^{\prime},q_{2}^{\prime})))=\lambda_{\mathcal{N}_{1}}((q_{1},q_{1}^{\prime}))\cap\lambda_{\mathcal{N}_{2}}((q_{2},q_{2}^{\prime}))$;
  • $Q_{S}=Q_{S1}\times Q_{S2}$;
  • $Q_{F}=Q_{F1}\times Q_{F2}$.
The resulting 2-NFA accepts the intersection of the languages defined by the initial automata $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$:
$[\![\mathcal{N}_{1}\times\mathcal{N}_{2}]\!]=[\![\mathcal{N}_{1}]\!]\cap[\![\mathcal{N}_{2}]\!].$
We have all the required preliminaries to formally state the problem solved by the 2-RPQ algorithm.
Definition 2.0 (Two-way regular path query).
Recall two-way regular path query (2-RPQ) a triple $\langle\mathcal{G},\mathcal{R},v_{s}\rangle$ where $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ is a graph, $\mathcal{R}$ is a regular language over the alphabet $L^{\leftrightarrow}$ and $v_{s}\in V$ is a starting node.
Frequent practical cases are queries with fixed starting or final vertex. So, for 2-RPQ $\mathcal{Q}=\langle\mathcal{G},\mathcal{R},v_{s}\rangle$ we are interested in an efficient way of evaluating these maps:
  • Single-source reachability:
    $[\![\mathcal{Q}]\!]_{SSR}=\left\{\;u\in V\;\middle|\;\begin{gathered}\exists\text{ 2-path }\pi\text{ in }\mathcal{G}\text{ from }v_{s}\text{ to }u\\ \text{and }\omega^{\leftrightarrow}_{\mathcal{G}}(\pi)\cap\mathcal{R}\neq\operatorname{\varnothing}\\ \end{gathered}\;\right\}$
  • Single-destination reachability:
    $[\![\mathcal{Q}]\!]_{SDR}=\left\{\;u\in V\;\middle|\;\begin{gathered}\exists\text{ 2-path }\pi\text{ in }\mathcal{G}\text{ from }u\text{ to }v_{s}\\ \text{and }\omega^{\leftrightarrow}_{\mathcal{G}}(\pi)\cap\mathcal{R}\neq\operatorname{\varnothing}\\ \end{gathered}\;\right\}$
3. Linear Algebra, Graphs and Relations
In order to operate with graphs in terms of linear algebra we need to see how algebraic object are connected with set relations and how the relations are connected to graph theory.
For two enumerated finite sets $A=\{1,2,...n\}$, $B=\{1,2,...,m\}$ for some $n,m\in\mathbf{N}$ a binary matrix $T$ of size $|A|\times|B|$ can be used to represent a relation $\mathcal{T}\subseteq A\times B$: $T_{ij}=1\Leftrightarrow(i,j)\in\mathcal{T}$. Recall this matrix $T$ a matrix representing the relation $\mathcal{T}$.
We need to define the following linear algebra operations over the Boolean matrices. Let $A$, $B$, $C$ be the matrices over Boolean algebra =$\langle\mathcal{B}=\{0,1\},\vee,\wedge,\neg,0,1\rangle$. Introduce the following operations and constants.
  • Sum: $A_{n\times m}\operatorname{\operatorname{\oplus}}B_{n\times m}=(a_{ij}\vee b_{ij})_{n\times m}$.
  • Product: $A_{n\times m}\operatorname{\operatorname{\otimes}}B_{m\times k}=\displaystyle\left(\;\bigvee_{k=1}^{n}a_{ik}\wedge b_{kj}\;\right)_{n\times k}$.
  • Transposition: $A_{n\times m}^{T}=(a_{ji})_{m\times n}$.
  • Zero-matrix: $\operatorname{\operatorname{\mathbf{0}}}_{n\times k}=(0)_{n\times k}$.
  • Mask: $A_{n\times m}\langle C_{n\times m}\rangle=(a_{ij}\wedge c_{ij})_{n\times m}$
  • Complement: $\neg A=(\neg a_{ij})$
Let $\mathcal{A}$, $\mathcal{B}$ and $\mathcal{C}$ be binary relations over sets $S_{1}$, $S_{2}$, $S_{3}$. Let $A$, $B$ and $C$ be binary matrices representing them. Then the following correspondences hold.
  • $C=A\operatorname{\operatorname{\oplus}}B$ $\leftrightarrow\mathcal{C}=\mathcal{A}\cup\mathcal{B}$
           where $\mathcal{A},\mathcal{B},\mathcal{C}\subseteq S_{1}\times S_{2}$.
  • $C=A\operatorname{\operatorname{\otimes}}B$ $\leftrightarrow\mathcal{C}=\{(i,k)\mid(i,j)\in\mathcal{A}$, $(j,k)\in\mathcal{B}\}$
           if $\mathcal{A}\subseteq S_{1}\times S_{2}$, $\mathcal{B}\subseteq S_{2}\times S_{3}$ and
           $\mathcal{C}\subseteq S_{1}\times S_{3}$.
  • $C=A^{T}$ $\leftrightarrow\mathcal{C}=\{(i,j)\mid(j,i)\in\mathcal{A}\}=\mathcal{A}^{-}$
           where $\mathcal{A}\subseteq S_{1}\times S_{2}$, $C\subseteq S_{2}\times S_{1}$.
  • $C=A\langle B\rangle$ $\leftrightarrow\mathcal{C}=\mathcal{A}\cap\mathcal{B}$,
           where $\mathcal{A},\mathcal{B},\mathcal{C}\subseteq S_{1}\times S_{2}$.
  • $C=\neg A$ $\leftrightarrow\mathcal{C}=\{(i,j)\mid(i,j)\in S_{1}\times S_{2}\setminus\mathcal{A}\}$
           where $\mathcal{A},\mathcal{C}\subseteq S_{1}\times S_{2}$.
Let $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ be an arbitrary graph and $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$ an arbitrary 2-NFA. Introduce the sets:
  • $E^{a}=\{e\mid e\in E,a\in\lambda^{\leftrightarrow}_{\mathcal{G}}(e)\}$ for all $a\in L^{\leftrightarrow}$.
  • $\Delta^{a}=\{\delta\mid\delta\in\Delta,a\in\lambda_{\mathcal{N}}(\delta)\}$ for all $a\in\Sigma^{\leftrightarrow}$.
These sets consist of the edges of the graph $\mathcal{G}$ and the transitions of 2-NFA $\mathcal{N}$ marked with a label $a$. They can be seen as binary relations over the sets $V$ and $Q$ correspondingly.
Definition 3.0 (Adjacency matrix of the label).
Let $G^{a}$ be the matrix representing the binary relation $E^{a}$. $G^{a}$ is called an adjacency matrix of the label $a$.
Definition 3.0 (Boolean decomposition of the adjacency matrix).
A Boolean decomposition of the adjacency matrix $G$ is a set of Boolean matrices $\{G^{a}\mid a\in L\}$.
Boolean decomposition of the adjacency matrices of some real-world data represented by graphs is a set of sparse matrices. This fact strictly leads to the idea of exploiting it for an efficient representation and using sparse matrix operation algorithm implementations.
Also note that for the given graph $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$, $E^{a^{-}}=(E^{a})^{-}$ and $G^{a^{-}}=(G^{a})^{T}$.
4. Single-Source RPQ in Terms of Linear Algebra
In this section we describe a single-source linear-algebra based 2-RPQ algorithm and prove its correctness.
input : $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}}Q_{S},Q_{F}\rangle$, $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$, $v_{s}\in V$
output : Matrix $P_{|Q|\times|V|}$
1
2$\{N^{a}\}\leftarrow$ Boolean decomposition of the $\mathcal{N}$ adjacency matrix;
3 $\{G^{a}\}\leftarrow$ Boolean decomposition of the $\mathcal{G}$ adjacency matrix;
4 $M_{|Q|\times|V|},P_{|Q|\times|V|}\leftarrow\operatorname{\operatorname{\mathbf{0}}}_{|Q|\times|V|}$ ;
5 for $q\in Q_{S}$ do
       $M_{q{v_{s}}}\leftarrow 1$;
        // Set the starting relation
6
7 end for
8while $M\neq\operatorname{\operatorname{\mathbf{0}}}$ do
       $M\leftarrow\bigoplus_{a\in\Sigma^{\leftrightarrow}\cap L^{\leftrightarrow}}((N^{a})^{T}\operatorname{\operatorname{\otimes}}M\operatorname{\operatorname{\otimes}}G^{a})\langle\neg P\rangle$;
        // Update $\mathcal{M}$
       $P\leftarrow P\operatorname{\operatorname{\oplus}}M$ ;
        // Accumulate resulting matrix
9
10 end while
11
return $P$
Algorithm 1 Single Source 2-RPQ using linear algebra
Let $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$ be the input 2-NFA represents query and specifies the RL $\mathcal{R}=[\![\mathcal{N}]\!]$. Let $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ be the input graph and $v_{s}\in V$ is a starting node. Then the algorithm 1 builds the following automata intersection:
$[\![\mathcal{N}\times\mathcal{N}_{G}]\!]=[\![\mathcal{N}]\!]\cap[\![\mathcal{N}_{\mathcal{G}}]\!]=\mathcal{R}\cap\left(\bigcup_{\begin{subarray}{c}\text{2-path }\pi_{\mathcal{G}}\text{ in }\mathcal{G}\\ \text{ from }v_{S}\in V_{S}\end{subarray}}\omega_{\mathcal{G}}^{\leftrightarrow}(\pi_{\mathcal{G}})\right).$
Core of the algorithm is a line 8 that performs one step of traversing two automata simultaneously. Result of the algorithm is a matrix $P_{|Q|\times|V|}$ such that
$\begin{cases}P_{q_{F}v}=1\\ q_{F}\in Q_{F}\end{cases}\Leftrightarrow\begin{cases}\exists\pi_{\mathcal{G}}\text{ in }\mathcal{G}\text{ from }v_{s}\text{ to }v\\ \omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\mathcal{R}\neq\operatorname{\varnothing}.\end{cases}$
An example of such a step is presented in figure 1. For graph and automaton with labels $\{a,b\}$, we visualise adjacency matrices for both of these symbols, the matrix of relation $M$, and the process of new $M$ computation. Solid-colored edges should be traversed simultaneously at the step. Dashed edges used to represent relation $M$: the edge connects two vertex that are simultaneously reachable. Initially, such an edge connects the initial state of the automaton and the starting vertex in the graph. After the step (one iteration of the main cycle), there are two such edges. One of them connects the final state of the automaton with vertex 5 of the graph. That means vertex 5 is reachable from the start vertex 3 by the path that farms a word acceptable by the automaton.
Hence, the set of reachable nodes can be easily obtained from the matrix $P$. Let $\mathcal{Q}=\langle\mathcal{G},\mathcal{R},v_{s}\rangle$ be an evaluated 2-RPQ. Then by the main property of the algorithm 1 resulting matrix $P$ is such that
$v\in[\![\mathcal{Q}]\!]_{SSR}\Leftrightarrow P_{q_{F}v}=1\text{ for some }q_{F}\in Q_{F}.$
Procedure 2 computes the explicit representation of the query result. Namely, $v\in[\![\mathcal{Q}]\!]_{SSR}\Leftrightarrow P^{F}_{1v}=1$ for some 2-RPQ $\mathcal{Q}=\langle\mathcal{G},\mathcal{R},v_{s}\rangle$. Formally, the following theorem holds.
Theorem 4.1 (LA 2-RPQ algorithm correctness).
The proposed algorithm, represented in 1, computes matrix $P$ such that respective relation $\mathcal{P}\subseteq Q\times V$ has the following property.
$(q,v)\in\mathcal{P}\Leftrightarrow\begin{cases}\exists\text{ 2-path }\pi_{G}\text{ in }\mathcal{G}\text{ from }v_{s}\text{ to }v\\ \exists\text{ path }\pi_{N}\text{ in }\mathcal{N}\text{ from some }q_{s}\in Q_{S}\text{ to }q\\ \omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}})\neq\operatorname{\varnothing}.\end{cases}$
input : Matrix $P_{|Q|\times|V|}$, set of final states $Q_{F}$
output : Matrix $P^{F}_{1\times|V|}$
1
[…]
3 for $q\in Q_{F}$ do
       $F_{1q_{F}}\leftarrow 1$;
        // Setting up a matrix reducing $P$
4
5 end for
return $P^{F}=F\operatorname{\operatorname{\otimes}}P$
Algorithm 2 Exporting reachable vertices subprogram
Proof of this theorem is straightforward induction by the paths length. Details are provided in appendix A.
Particularly, from the theorem, we can immediately conclude that for the given automaton $\mathcal{N}$, graph $\mathcal{G}$ and starting vertex $v_{s}$, $P_{q_{F}v}=1$ for $q_{F}\in Q_{F}$ iff exist two paths $\pi_{\mathcal{G}}$ to $v$ and $\pi_{\mathcal{N}}$ to $q_{F}$ such that $\omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}})\neq\operatorname{\varnothing}$. As far as $q_{F}$ is a final state of the NFA, $\omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\subseteq[\![\mathcal{N}]\!]$. Thus, $v$ reachable from $v_{s}$ by path satisfies the given regular constraint.
Figure 1. The algorithm step for graph and automaton for regular expression $a^{*}b$
The proposed single-source 2-RPQ algorithm can be used to solve the single-destination problem. Let $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$ be an NFA specifies the constraint in single destination query. Denote:
  • =$\mathcal{R}^{-}=\{w^{-}=a_{n}^{-}a_{n-1}^{-}...a_{2}^{-}a_{1}^{-}\mid w=a_{1}a_{2}...a_{n})\in\mathcal{R}\}$
  • $\mathcal{N}^{-}=\langle Q,\Sigma,\Delta^{-},\lambda_{\mathcal{N}}^{-},Q_{F},Q_{S}\rangle$.
  • $\mathcal{Q}^{-}=\langle\mathcal{G},\mathcal{R}^{-},v_{s}\rangle$.
The automaton $\mathcal{N}^{-}$ allows traversing paths in the opposite direction. It holds $\mathcal{R}^{-}=[\![\mathcal{N}^{-}]\!]$. Evaluating single-source reachability 2-RPQ on the reversed 2-RPQ $\mathcal{Q}^{-}$ gives using the 2-NFA $\mathcal{N}^{-}$:
$[\![\mathcal{Q}]\!]_{SDR}=[\![\mathcal{Q}^{-}]\!]_{SSR}.$
Boolean matrix decomposition of $\mathcal{N}^{-}$ can be easily obtained by transposing the matrices in Boolean matrix decomposition of $\mathcal{N}$.
For performance reasons, it’s necessary to take into account the fact that sparse matrix multiplication works faster when matrices have fewer non-zero entries. Associativity of the Boolean matrix multiplication can be used to perform less calculation. This leads to the idea of two possible ways to calculate the product on line 8 in the algorithm 1:
$((N^{a})^{T}\operatorname{\operatorname{\otimes}}M)\operatorname{\operatorname{\otimes}}G^{a}=(N^{a})^{T}\operatorname{\operatorname{\otimes}}(M\operatorname{\operatorname{\otimes}}G^{a}).$
In general, the first option is preferable due to the fact that 2-NFA for the query can be converted into the corresponding minimal deterministic finite automaton having no more than $|Q|$ non-zero cells in each $\mathcal{N}^{a}$ adjacency matrix due to the automaton determinism. This strictly leads to less expensive computation of the product with large $|V|\times|V|$ matrix.
5. Evaluation
The proposed single-source 2-RPQ algorithm is implemented using the SuiteSparse:GraphBLAS library (Davis, 2019b) inside the LAGraph (Szárnyas et al., 2021) infrastructure [^6] The source code of the algorithm is available in our GitHub LAGraph fork: https://github.com/SparseLinearAlgebra/LAGraph/tree/rpq. Both sparse matrices and their transpositions are loaded in-memory due to the fact we need both representations to efficiently evaluate 2-RPQ and traverse the graph in both directions.
Whereas we’re focusing on an efficient way of evaluating RPQs in-memory, there aren’t many candidates to compare with. Popular database management systems primarily focus on general availability and ensure availability of the concurrent access at the same time the suggested algorithm implementation is suited only for solving the reachability problem. We’ve selected the following systems as demonstrating the best performance and being the most related to our use case.
MillenniumDB (Vrgoc et al., 2021) is a graph-oriented database management system supporting a synthetic way to carry out calculations without using the disk storage by caching the query data. It demonstrates state-of-the-art performance on evaluating regular path queries.
FalkorDB (ex. RedisGraph (Cailliau et al., 2019)) is an in-memory property-graph database also employing SuiteSparse:GraphBLAS for query evaluation. However, it uses OpenCypher modification and supports only a subset of regular path queries. To deal with this, we’ve carried out the measurements of cypher-compatible and non-cypher-compatible queries separately (e.g., it is impossible to evaluate repeated-path queries in the form of $(ab)^{*}$ in FalkorDB).
RPQ-matrix (Arroyuelo et al., 2023a) is another implementation of the in-memory linear-algebra based RPQ evaluation algorithm. The original work introduces a few variations of the adjacency matrix representation: $k^{2}$-trees offering less memory consumption and CSR/CSC formats providing better performance. We’ve chosen the last one to compare with since we’re aiming to compare the performance.
All experiments are conducted on a work station with Ryzen 9 7900X 4.7 GHz 12-core, 128 Gb of DDR5 RAM and running Ubuntu 22.04.
5.1. Dataset Description
For algorithm evaluation, we choose the Wikidata dataset from the snapshot provided in terms of MillenniumDB path query challenge (Farias et al., 2023). The resource contains both the graph and the set of 660 different 2-RPQs in SPARQL format taken from the Wikidata query log.
For systems that do not support RDF format, the dataset was converted to an edge-labelled graph. The processed graph has 610 402 396610402396610\,402\,396610 402 396 edges, 91 609 2549160925491\,609\,25491 609 254 vertices and 1 39513951\,3951 395 distinct labels.
SPARQL path queries has been converted to the corresponding minimized DFAs. For evaluating on FalkorDB the supported subset of regular path queries have been converted into Cypher queries of form MATCH ... COUNT (DISTINCT ...). The original dataset contains queries from all to all vertices and queries involving non-existing entities. They’ve been filtered out, resulting in 578 queries remaining.
5.2. Evaluation Results
Figure 2. Wikidata dataset per-query evaluation time
LA 2-RPQ MilenniumDB RPQ-Matrix FalkorDB
Mean C (ms) 83 327 348 1 625
Median C (ms) 2.5 4.4 4.2 18.8
Mean speedup C 1 3.9 4.2 19.6
Median speedup C 1 1.8 1.7 7.5
Mean NC (ms) 374 1 345 7 108
Median NC (ms) 37.4 69.8 269.6
Mean speedup NC 1 3.6 19.0
Median speedup NC 1 1.9 7.2
Memory (Gb) 18.1 6.9 50.0
BPT 31.9 12.1 85.8
Timeouts 0 1 15 53
Table 1. Query execution results. Results are provided for cypher-compatible (C) and non-cypher-compatible (NC) queries separately. Mean and median speedup relatively the proposed solution is a ratio of mean and median over query set for respective systems
We measure evaluation time per query with a timeout of 1 minute. All of the data required for query evaluation is pre-loaded into the memory. Queries are executed one by one.
We run MillenniumDB [^7] MillenniumDB repository we use for the evaluation: https://github.com/MillenniumDB/MillenniumDB/tree/5190c0d9b07ca681328495b69c715af792513775 in a SPARQL mode with the maximum amount of CPU cores available. The query dataset is executed twice. The first run isn’t measured since it makes the database load the data into the memory. The query execution time is calculated as the internal optimizer plus execution time excluding the parser.
We execute FalkorDB v4.2.0888FalkorDB v4.2.0: https://github.com/FalkorDB/FalkorDB/tree/v4.2.0 querying through the python wrapper. The database is started without CPU core limitations and without query cache. We build an index over the vertices to be able to select the starting or destination vertex using it. The query execution time is based on the internal timer.
The RPQ-matrix algorithm is executed in CSR/CSC mode without core and memory consumption limitations. The consumed time is measured with the internal timer. The solution has been built out of the project repository [^9] RPQ-Matrix repository we use for the evaluation: https://github.com/adriangbrandon/rpq-matrix/tree/34fc2240a7c8069f7d6a39f1c75176edac4fe606.
Results of the measurements, both evaluation time and relative speedup, are provided in figure 2. Only queries expressible in Cypher are presented. The purple lines represent mean values, and the green lines represent medians. Note that relative speedup is represented as pointwise speedup for each query. So, mean and median calculated for per-query speedup. Numeric values of the mean and the median, the size of the dataset loaded in-memory, and memory byte-per-triple (BPT) memory consumption values are available within the table 1. Queries that have failed on any of the systems with timeouts are excluded from the statistics. The number of time outed queries shown in table 1.
We can see that while in median, speeding up for our solution is between 1.7 and 7.5, in mean we improve performance more than 19 times (relative to the most closest solutions RPQ-Matrix and FalkorDB). That means the proposed solution significantly outperforms other linear algebra based solutions on some queries (up to ×$\times 10^{5}$) while demonstrating comparable (and better in average) performance in the most cases. On the other hand, RPQ-Matrix demonstrates better performance on big number of queries. It is an interesting question for future research whether it is possible to combine ideas from two linear algebra based solutions (proposed and RPQ-Matrix) to take the best from both of them.
Note that MillenniumDB is pretty stable: there is neither outstanding wins nor losses in performance. At the same time, the proposed solution is slightly better in average even though MillenniumDB caches query evaluation data, but our solution does not.
Memory consumption is presented on table 1. We don’t specify MillenniumDB memory consumption due to the fact it can process query in-memory only by caching data and doesn’t provide ways to preload whole dataset from the disk storage. RPQ-Matrix demonstrates the lowest memory usage. Memory consumption of the proposed solution is comparable with competitors.
We collected the top 10 queries with the worst and best speedup relative to all competitors. Collected data presented in tables 345. We can see that the weakness of our solution are relatively simple queries with a relatively small number of reachable vertices, like Q P131* ?x1 with less than 100 reachable vertices, or ?x1 (P138 / P106) Q with less than a million reachable vertices, where Q is a fixed vertex, and ?x1 is a variable. The forte of the proposed solution is a complex query (in terms of regexp complexity) with a huge number of reachable vertices.
6. Conclusion and Future Work
In this work, we proposed the single-source RPQ evaluation algorithm, which is based on simultaneous traversal of the input graph and the finite automaton specifying path constraints. The traversal is expressed in terms of operations over matrices and vectors, that allows us to provide highly-parallel implementation based on SuiteSparse:GraphBLAS. Our evaluation shows that the suggested algorithm is suitable for in-memory processing of real-world large knowledge graphs. While for some queries the proposed algorithm is slower than competitors, we can conclude that our solution is faster for hard queries: it fits with 1 minute time while other solutions do not.
First of all, it is necessary to analyse abilities to apply well-known optimizations from both RPQ evaluation and BFS algorithms. For example, rare labels (Koschmieder and Leser, 2012) utilization, or push-pull optimization (Yang et al., 2022) respectively.
While it was shown that multiple source BFS can be expressed in terms of linear algebra (Elekes et al., 2020), there is a room for technical optimizations and careful evaluation of the respective modifications to the proposed algorithm, and it should be done in the future.
Thanks to linear algebra, having a single schema of algorithm one can solve different problems varying semiring-like structures. For example, one can look at variations of BFS (Brock et al., 2021) where one can compute reachability facts or information about paths depending on used semiring. In the case of RPQ, there are a number of possible outputs and semantics (Angles et al., 2017): reachability, single path, all paths, simple paths, etc. It is an open question, which of them can be expressed without algorithm changes, but by providing other semiring, and which can be expressed with algorithm changes.
Utilization of GPGPUs to evaluate linear algebra based algorithms for graph analysis can significantly improve performance in some cases (Yang et al., 2022; Stanislavovic, 2023). It is necessary to investigate, whether utilization of GPGPU in our algorithm improve performance or not.
Distributed solutions are a way for huge graphs processing (Guo et al., 2021b). Implementation and evaluation of the proposed algorithm in distributed settings, for example, using CombBLAS (Buluç and Gilbert, 2011) that provides distributed linear algebraic routines for graph analysis, is also a task for the future.
Acknowledgements.
This research was supported by the St Petersburg University, grant id 116636233.
Appendix A Proof of Correctness
Theorem A.1 (LA 2-RPQ algorithm correctness).
The proposed algorithm, represented in 1, computes matrix $P$ such that respective relation $\mathcal{P}\subseteq Q\times V$ has the following property.
$(q,v)\in\mathcal{P}\Leftrightarrow\begin{cases}\exists\text{ 2-path }\pi_{G}\text{ in }\mathcal{G}\text{ from }v_{s}\text{ to }v\\ \exists\text{ path }\pi_{N}\text{ in }\mathcal{N}\text{ from some }q_{s}\in Q_{S}\text{ to }q\\ \omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}})\neq\operatorname{\varnothing}.\end{cases}$
Label Count
P31 95 749 143
P17 13 994 387
P131 10 609 188
P106 8 533 132
P279 3 128 156
P19 2 800 454
P40 1 419 672
P39 1 302 690
P495 1 272 895
P20 1 127 738
P138 322 330
P551 237 546
Table 2. Number of edges with specific labels used in queries
Query Speedup #Answer
1 Q13442814 ^P31 ?x1 19.09 37 280 216
2 ?x1 (P17 / P31) Q6256 14.49 12 360 733
3 ?x1 (P31* / P279*)+ Q268592 13.34 60 480 400
4 ?x1 ((P106 | P39 | P31) / P279*) Q488383 12.79 76 388 701
5 ?x1 (P31 / P279*) Q191067 11.72 38 410 384
6 ?x1 (P31 / P279*) Q488383 11.10 75 238 748
7 ?x1 (P31 / P279*) Q35120 10.50 90 321 703
8 ?x1 (P31 / P279*) Q13442814 10.44 37 325 506
9 ?x1 (P31 / P279*) Q386724 9.71 55 722 209
10 ?x1 (P31 / P279*) Q7184903 9.53 58 080 849
11 Q10281806 P131* ?x1 0.16 6
12 Q1628288 P131* ?x1 0.17 6
13 Q111 (P31 / P279*) ?x1 0.19 18
14 Q939035 P131* ?x1 0.19 6
15 Q1549583 P131* ?x1 0.20 6
16 Q481075 P131* ?x1 0.20 5
17 Q29873614 P131* ?x1 0.21 4
18 ?x1 P40+ Q748955 0.21 17 699
19 Q28733811 P131* ?x1 0.21 5
20 ?x1 P40+ Q390192 0.23 22 335
Table 3. Top 10 relatively slowest and relatively fastest queries for MDB
Denote state of matrix $M$ after the step $n$ of the loop on lines 7–10 of the algorithm 1 as $M_{n}$ and state of $P$ as $P_{n}$. Introduce auxiliary relations $\mathcal{M}_{n}\subseteq Q\times V$ represented by $M_{n}$ and $\mathcal{P}_{n}\subseteq Q\times V$ represented by $P$ after the step $n$. $\mathcal{P}$ and $\mathcal{P}_{n}$ are connected with relations $\mathcal{M}_{n}$:
$\mathcal{P}_{n}=\bigcup_{1\leq m\leq n}\mathcal{M}_{m};\;\;\mathcal{P}=\bigcup_{m\in\mathbb{N}}\mathcal{M}_{m}$
The approach is to build a relation between the states of automaton and the vertices of the graph and update it by traversing both automaton and the graph at the same time. As an algorithm invariant after the step $n$ we claim this properties for $\mathcal{M}_{n}$ represented by the matrix $M_{n}$:
$(q,v)\in\mathcal{M}_{n}\Leftrightarrow\begin{cases}\exists\text{ 2-path }\pi_{\mathcal{G}}\text{ of length }n\text{ in }\mathcal{G}\text{ from }v_{s}\text{ to }v\\ \exists\text{ path }\pi_{\mathcal{N}}\text{ of length }n\text{ in }\mathcal{N}\text{ from }q_{s}\in Q_{S}\text{ to }q\\ \omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}})\neq\operatorname{\varnothing}\\ \forall m<n\;(q,v)\notin\mathcal{M}_{m}\end{cases}$
Note, that if we continue to loop forever in lines 7–10 from $\mathcal{M}_{n}=\operatorname{\varnothing}$ it follows $\mathcal{M}_{m}=\operatorname{\varnothing}$ for $m>n$. It means, it’s enough to iterate until the $M$ matrix becomes $\operatorname{\operatorname{\mathbf{0}}}$.
Notice, that $\mathcal{P}_{n}\subsetneq\mathcal{P}_{n+1}$ or $|\mathcal{P}_{n}|<|\mathcal{P}_{n+1}|$ if $\mathcal{M}_{n+1}$ is not empty. And $\mathcal{M}_{n+1}\cap\mathcal{P}_{n}=\operatorname{\varnothing}$ for all $n\in\mathbb{N}$. $|\mathcal{P}_{n}|\leq|Q||V|$ and $|\mathcal{M}_{n}|\leq|Q||V|$ for all $n\in\mathbb{N}$. Thus, if $\mathcal{M}_{m}\neq\operatorname{\varnothing}$ for all $m<|Q||V|$ then $\mathcal{M}_{|Q||V|}=\operatorname{\varnothing}$, since $|\mathcal{P}_{|Q||V|-1}|\geq|Q||V|$. This means the algorithm always finishes in maximum of $|Q||V|$ steps.
Obviously, the invariant is held for $n=0$ after the initialization of the matrix $M$ with $M_{0}$. If you consider the paths of length 00 the set of vertices and automaton states coincides with $\{v_{s}\}$ and $Q_{S}$ correspondingly.
$(q,v)\in\mathcal{M}_{0}\Leftrightarrow q\in Q_{S},v=v_{s}$
Consider evaluating the $n+1$ step of the algorithm. Fix the label $a\in L$. After the step $n$ $M$ represents a relation $\mathcal{M}_{n}$. At first, evaluate the first matrix product $M^{\prime a}_{n+1}=(N^{a})^{T}\operatorname{\operatorname{\otimes}}M$. This product represents a relation $\mathcal{M}^{\prime a}_{n+1}\subseteq Q\times V$:
$\mathcal{M}^{\prime a}_{n+1}=\{(q^{\prime},v)\mid(q,q^{\prime})\in\Delta^{a},(q,v)\in\mathcal{M}_{n}\}.$
Evaluate the second matrix product $M^{a}_{n+1}=M^{\prime a}_{n+1}\operatorname{\operatorname{\otimes}}G^{a}=$
$=(N^{a})^{T}\operatorname{\operatorname{\otimes}}M_{n}\operatorname{\operatorname{\otimes}}G^{a}$. Assume $M^{a}_{n+1}$ represents a relation $\mathcal{M}^{a}_{n+1}$, then:
$\mathcal{M}_{n+1}^{a}$ =$=\{(q^{\prime},v^{\prime})\mid(v,v^{\prime})\in E^{a},(q^{\prime},v)\in\mathcal{M}^{\prime a}_{n+1}\}$
=$=\{(q^{\prime},v^{\prime})\mid(v,v^{\prime})\in E^{a},(q,q^{\prime})\in\Delta^{a},(q,v)\in\mathcal{M}_{n}\}.$
Hence, the new relation $\mathcal{M}_{n+1}$ represented by the matrix
$\bigvee_{a\in\Sigma^{\leftrightarrow}\cap L^{\leftrightarrow}}M^{a}_{n+1}\langle\neg P\rangle$ can be written as follows:
$\mathcal{M}_{n+1}$ $=\bigcup_{a\in\Sigma^{\leftrightarrow}\cap L^{\leftrightarrow}}\{(q,v)\mid(q,v)\in\mathcal{M}^{\prime a}_{n+1},(q,v)\notin\mathcal{P}\}=$
$=\bigcup_{a\in\Sigma^{\leftrightarrow}\cap L^{\leftrightarrow}}\left\{(q^{\prime},v^{\prime})\middle|\begin{gathered}(v,v^{\prime})\in E^{a},(q,q^{\prime})\in\Delta^{a}\\ (q,v)\in\mathcal{M}_{n},(q,v)\notin\mathcal{P}\end{gathered}\right\}.$
The updating the relation between the 2-NFA states and the graph vertices $\mathcal{M}$ with $\mathcal{M}^{\prime}$ value is follows form the following properties.
$(q^{\prime},v^{\prime})\in\mathcal{M}_{n+1}\Leftrightarrow\begin{cases}(q,q^{\prime})\in\Delta^{a}\\ (v,v^{\prime})\in E^{a}\\ \end{cases}\text{ for some }a\in\Sigma^{\leftrightarrow}\cap L^{\leftrightarrow}.$
Ensure the invariant is preserved for $(q^{\prime},v^{\prime})$ in $\mathcal{M}^{\prime}$. Let $\pi_{\mathcal{G}}=(e_{1},...,e_{n})$, $\pi_{\mathcal{N}}=(\delta_{1},...,\delta_{n})$ the paths to $(q,v)$ satisfying the conditions A. Then:
  • $\exists$ 2-path $\pi_{\mathcal{G}}^{\prime}=(e_{1},...,e_{n},(v,v^{\prime}))$ in $\mathcal{G}$ from $v_{s}$ to $v^{\prime}$.
  • $\exists$ path $\pi_{\mathcal{N}}^{\prime}=(\delta_{1},...,\delta_{n},(q,q^{\prime}))$ in $\mathcal{N}$ from $q_{F}\in Q_{F}$ to $q^{\prime}$.
  • $w\cdot a\in\omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}}^{\prime})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}}^{\prime})$ where $w\in\omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}})$.
  • $\mathcal{P}_{n}=\displaystyle\bigcup_{m\leq n}\mathcal{M}_{m}$, thus, if $(q,v)\in\mathcal{M}_{m}$ for $m\leq n$ then $(q,v)\notin\mathcal{M}_{n+1}$.
Since, the conditions (A) are preserved for the relation $\mathcal{M}_{n+1}$ matching the new value of $M\leftarrow M_{n+1}$ the invariant is preserved.
Appendix B Queries Statistics
For detailed analysis of performance we collected top 10 relatively fastest and relatively slowest queries for each system. Results presented in tables 345. Queries represented in SPARQL-like syntax. Speedup means speeding up of the proposed solution relative to respective system. #Answer is a number of reachable vertices. Also we collected total number of edges in the graph with specific labels appeared in collected queries. Results presented in tables 2.
Query Speedup #Answer
1 Q7187 ((P279 / (P279* | P31*)) | (P31/(P279* | P31*))) ?x1 506 771.72 18
2 Q5 ((P279 / (P279* | P31*)) | (P31 / (P279* | P31*))) ?x1 345 401.17 30
3 Q11629 ((P279 / (P279* | P31*)) | (P31 / (P279* | P31*))) ?x1 311 512.45 63
4 Q1656682 ((P279 | P31) / (P279* | P31*)) ?x1 306 341.47 7
5 Q5 (P279 / (P279* | P31*)) ?x1 241 437.89 20
6 Q5 ((P279 | P31) / (P279* | P31*)) ?x1 185 795.87 30
7 Q81328 (P31 / P279? / P279? / P279? / P279? / P279?) ?x1 127 813.38 18
8 Q4235 (P31 / P279? / P279? / P279? / P279? / P279?) ?x1 123 845.99 18
9 Q2 (P31/P279? / P279? / P279? / P279? / P279? / P279?) ?x1 122 229.53 20
10 Q111 (P31 / P279? / P279? / P279? / P279? / P279? / P279?) ?x1 92 952.34 11
11 ?x1 (P17 | P495) Q30 0.03 1 697 681
12 ?x1 (P138 / P106) Q1028181 0.04 8 135
13 ?x1 (P31 / P279) Q18918145 0.05 188 351
14 ?x1 (P19 | P20 | P551) Q90 0.06 50 972
15 ?x1 (P31 / P279) Q16521 0.07 245 925
16 ?x1 (P19 | P20 | P551) Q1741 0.07 43 405
17 ?x1 (P31 / P279*) Q18918145 0.09 217 394
18 ?x1 P31+ Q4830453 0.09 190 623
19 ?x1 (P31 / P279) Q1030034 0.09 82 052
20 ?x1 (P31 / P279) Q486972 0.09 581 652
Table 4. Top 10 relatively slowest and relatively fastest queries for RPQ-Matrix
Query Speedup #Answer
1 ?x1 (P31 / P279*) Q959309 37327.83 1605
2 ?x1 (P31 / P279*) Q385337 30592.48 1980
3 ?x1 (P31/P279*) Q543654 10783.13 7318
4 Q64 (P31/P279*) Q811430 8052.45 115
5 ?x1 P279* Q16707842 2000.93 59
6 ?x1 P279* Q15630906 1664.48 13
7 ?x1 P131* Q168159 1335.28 267
8 ?x1 P279* Q144 606.69 269
9 ?x1 P131* Q108053 598.19 2137
10 ?x1 (P31 / P279*) Q1867183 596.26 1217
11 ?x1 (P138 / P106) Q1028181 0.66 8 135
12 Q257023 P131* ?x1 1.00 6
13 Q326313 P131* ?x1 1.19 6
14 Q204756 P131* ?x1 1.23 5
15 Q29018634 P131* ?x1 1.26 7
16 Q24192592 P131* ?x1 1.26 7
17 Q5964731 P131* ?x1 1.29 6
18 Q328046 P131* ?x1 1.30 6
19 Q2321874 P131* ?x1 1.33 5
20 Q894219 P131* ?x1 1.34 5
Table 5. Top 10 relatively slowest and relatively fastest queries for FalkorDB